OpenStudy (anonymous):

Two planets - planet A and planet B - are orbiting a star.If planet A has an orbital radius which is four times as large as planet B then the period of planet A orbit is _ times larger than the period of Planet B orbit

OpenStudy (jchick):

Applying Kepler's law of harmonies to this situation would result in: TA2 / RA3 = TB2 / RB3 This equation can be algebraically rearranged to TA2 / TB2 = RA3 / RB3 The ratio of the period squared of planet A to planet B will be equal to the ratio of the radius cubed of planet A to planet B. The ratio of the radii of the two planets is given - planet A's radius is two (or three) times larger than planet B's radius. The cube of this ratio is equal to the square of the ratio of the period. Taking the square root of the period squared ratio will yield the ratio of the periods of the planets. Mathematically, this could be written as TA / TB = SQRT(TA2 / TB2) = SQRT(RA3 / RB3)

OpenStudy (anonymous):

answer 8

OpenStudy (anonymous):

can u plug in the number

OpenStudy (anonymous):

can u help me plug in the numbers

OpenStudy (jchick):

@pooja195 I am blanking out on this can you help?

OpenStudy (jchick):

@Loser66 can you help I am blanking out.

pooja195 (pooja195):

@zepdrix

OpenStudy (jchick):

Ta2 / Tb2 = Ra3 / Rb3

OpenStudy (jchick):

Sorry I am trying to remember.

OpenStudy (anonymous):

i know the answer is 8 but i cant put it in the formula

OpenStudy (jchick):

Wait how do you know the answer is 8 if you don't know how to input it?

OpenStudy (anonymous):

i didnt use the formula thats why i tried doing 4^3=sqrt64 which equals 8

OpenStudy (jchick):

I can't remember sorry I wouldn't want to give you the wrong information.

OpenStudy (anonymous):

can u help me understand the formula

OpenStudy (jchick):

T2 = kr3 . In every day terms, T is the time of the orbital period of the planet, K is a constant, and r is the ratio of the Sun to the planet. If more than on planet is involved, the ratio becomes (Ta2 / Tb2 ) = (Ra3 / Rb3 ). This is still used today to find distances of objects found inside the Milky Way.

OpenStudy (anonymous):

can u plug it in

OpenStudy (jchick):

Can you tell me what eight is?

OpenStudy (anonymous):

eight is the answer

OpenStudy (jchick):

Ok answer to what. The whole problem? Because I cannot give you straight answers I want you to help come to the answer so that you understand it.

OpenStudy (jchick):

Where did you get the answer?

OpenStudy (anonymous):

i did this √4³ = 8

OpenStudy (anonymous):

this is the answer to the whole problem but i dont know how to put it into your formula

OpenStudy (jchick):

Are you working with Kepler?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

TA / TB = SQRT(TA2 / TB2) = SQRT(RA3 / RB3) is this the formula

OpenStudy (jchick):

Yes now do you know what the satellite mean orbital radius is?

OpenStudy (anonymous):

planet a orbital radius is 4 times as larger as planet B

OpenStudy (jchick):

Ok so what is planet B

OpenStudy (jchick):

What is the planetary mass?

OpenStudy (anonymous):

2 pie idk

OpenStudy (jchick):

Here is a file that will help

OpenStudy (anonymous):

By performing simple mathematics for planet A and B T(A)^2 / T(B)^2 = R(A)^3 / R(B)^3 Given R(A) = 4X R(B) T(A)^2 / T(B)^2 = 64 (As 4X4X4=64) T(A) / T(B) = 8 Thus Planet A takes 64 takes 8 times of Time Period over Planet B so answer is 8 times

OpenStudy (jchick):

Correct how did you find it?

OpenStudy (anonymous):

looked at the previous formula and assume the radius

OpenStudy (jchick):

Nice!

OpenStudy (jchick):

Did you see my file though?

OpenStudy (anonymous):

yes thanks for the help

OpenStudy (jchick):

No problem sorry I couldn't help more.